OpenStudy (anonymous):

Evaluate (sin 4x dx) on the interval from 0 to pi/4; what is the definite integral?

OpenStudy (zarkon):

\[\int\limits_{0}^{\pi/4}\sin(4x)dx=\frac{1}{2}\]